A dipole is consisting of an electron and proton, 4e-10 m apart. compute the electric field at a distance of 2e-8 m on a line making perpendicular with the dipole exist from the centre of the dipole?
Answer this question please
Thanks
1
Expert's answer
2019-12-11T10:30:03-0500


For the situation shown above, we can first calculate the distance from the charges to the considered point:


"d=\sqrt{(4\cdot10^{-10})^2+(2\cdot10^{-8})^2}=2\cdot10^{-8}\text{ m}."

Now calculate the magnitude of the vectors of the electric field lines at the considered point:


"E=E_-=E_+=ke\/d^2."

However, the field has two components: vertical and horizontal. The horizontal component of the total resultant field is


"E_x=E_+\text{ cos}\alpha+E_-\text{ cos}\alpha=2E\text{ cos}\alpha=\\
=179751036\text{ V/m}."

The vertical component of the resultant field is


"E_y=E_+\text{ sin}\alpha-E_-\text{ sin}\alpha=0."